OpenStudy (anonymous):

the line perpendicular to (3/2)y=(-4/7)x+3 and passing through (7,3)

OpenStudy (tyteen4a03):

Firstly find the slope of the line of (3/2)y=(-4/7)x+3. This is the slope-intercept form of the equation. We divide both sides by 3/2: \(y = \frac {-8} {21}x + 2\) The slope of this line is -8/21. Perpendicular lines have negative reciprocals of each other. In this case, the slope for the new line would be 21/8. Now plug the numbers in the point-slope form: y - 3 = (21/8)(x - 7) And there you have your answer.

OpenStudy (anonymous):

@tyteen4a03 thank you :)